# PFML: Self-Supervised Learning of Time-Series Data Without Representation Collapse

### v1

@@ Abstract
Self-supervised learning (SSL) is a data-driven learning approach that utilizes the innate structure of the data to guide the learning process. In contrast to supervised learning, which depends on external labels, SSL utilizes the inherent characteristics of the data to produce its own supervisory signal. However, one frequent issue with SSL methods is representation collapse, where the model outputs a constant input-invariant feature representation. This issue hinders the potential application of SSL methods to new data modalities, as trying to avoid representation collapse wastes researchers’ time and effort. This paper introduces a novel SSL algorithm for time-series data called Prediction of Functionals from Masked Latents (PFML). Instead of predicting masked input signals or their latent representations directly, PFML operates by predicting statistical functionals of the input signal corresponding to masked embeddings, given a sequence of unmasked embeddings. The algorithm is designed to avoid representation collapse, rendering it straightforwardly applicable to different time-series data domains, such as novel sensor modalities in clinical data. We demonstrate the effectiveness of PFML through complex, real-life classification tasks across three different data modalities: infant posture and movement classification from multi-sensor inertial measurement unit data, emotion recognition from speech data, and sleep stage classification from EEG data. The results show that PFML is superior to a conceptually similar pre-existing SSL method and competitive against the current state-of-the-art SSL method, while also being conceptually simpler and without suffering from representation collapse. The code is freely available at https://github.com/SPEECHCOG/PFML.
@@ 1 Introduction
    We show that PFML obtains both superior results against a conceptually similar pre-existing SSL method, and competitive results against the current state-of-the-art data modality agnostic SSL method, while also being conceptually simpler and without suffering from representation collapse.
@@ 2 Related Work
Similar to the advances in SSL for audio data, there have been significant developments in SSL for image data as well (Lee et al., 2017; Gidaris et al., 2018; Caron et al., 2018; Grill et al., 2020; Chen et al., 2020; Radford et al., 2021; He et al., 2022; Bao et al., 2022; Oquab et al., 2024). Grill et al. (2020) presented an SSL method that uses two neural networks that learn from each other’s representations of differently augmented views of the same image. He et al. (2022) proposed masked autoencoders (MAE) that try to reconstruct masked patches of input images using an asymmetric encoder-decoder architecture. The SSL algorithm by Bao et al. (2022) tokenizes images into visual tokens, followed by masking some image patches and then trying to recover the original tokens from the masked patches.
@@
SSL has also excelled in natural language processing (Devlin et al., 2019; Brown et al., 2020; Tay et al., 2023; OpenAI, 2023). Devlin et al. (2019) introduced an SSL method which obtains bidirectional feature representations from unlabeled text by conditioning on both the left and right textual context. The method by Brown et al. (2020) uses an autoregressive model which alternates dense and locally banded sparse attention patterns in their Transformer model. OpenAI (2023) proposed an expanded version of Brown et al. (2020) by making the model not only larger, but also capable of handling image inputs in addition to text inputs.
@@
More recently, SSL literature has seen a growing number of work towards SSL algorithms capable of running the pre-training task on multiple different data modalities. The authors of van den Oord et al. (2018) developed an SSL approach that predicts future embeddings based on previous context using contrastive learning. They showed that their method was able to learn useful feature representations for audio, image, text, and reinforcement learning in 3D environments. The SSL method by Akbari et al. (2021) also uses contrastive learning, but their method simultaneously takes audio, video, and text data as input and creates multimodal feature representations. These features were shown to work well with multiple different downstream tasks, i.e. video action recognition, audio event classification, image classification, and text-to-video retrieval. Wang et al. (2023) proposed an SSL method that performs prediction of masked tokens in a unified manner on images, texts, and image-text pairs. Their experiments showed that their method achieves state-of-the-art performance on various vision and vision-language tasks. Baevski et al. (2022) proposed data2vec, an SSL method for audio, image, and text data. In their approach, the model tries to predict masked latent features of an older version of itself that are both normalized and averaged over multiple Transformer layers. Their results in downstream tasks demonstrate the effectiveness of the method in all three data modalities.
@@
For modality agnostic SSL algorithms, objective functions play a crucial role in guiding the learning process. These functions can be broadly categorized into three types: instance discrimination, clustering, and masked prediction. Instance discrimination aims to distinguish between different instances of data, thereby encouraging the model to learn unique features for each instance and enhancing the discriminative power of the learned representations. Contrastive learning methods, such as van den Oord et al. (2018); Baevski et al. (2020); He et al. (2020); Akbari et al. (2021); Pizzi et al. (2022), are an example of instance discrimination-based SSL methods. Clustering, on the other hand, groups similar instances together in the feature space, fostering the model to learn common features among instances belonging to the same group. Methods like Caron et al. (2020); YM. et al. (2020); Hsu et al. (2021) are examples of clustering-based SSL methods. Lastly, masked prediction involves the task of predicting masked parts of the input data based on the unmasked parts, thereby encouraging the model to learn contextual relationships within the data. Examples of such SSL methods include Devlin et al. (2019); Wang et al. (2020); Baevski et al. (2022); He et al. (2022); Xie et al. (2022).
@@ 3.2 Prediction of Functionals from Masked Latents
Under these assumptions, as the model is trying to predict the computed functionals $\mathbf{f}_{n}$ given the embeddings $\mathbf{z}_{n}$, good model predictions $\mathbf{y}_{n}$ that lead to low prediction loss values also inherently contain variance. On the contrary, if $\mathbf{y}_{n}$ were to contain zero variance across the frames while $\mathbf{f}_{n}$ contains variance, the prediction loss would be high. Consequently, PFML pre-training does not converge to collapsed feature representations, as long as Assumptions 1 and 2 hold true. For a more detailed formulation, see Appendix A. Empirical results (see Section 4.4) support this theoretical claim, showing that PFML maintains variance in predictions across various datasets.
@@ 4 Experiments
We evaluate our PFML method using three different datasets of time-series data with complex classification tasks: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. For each dataset, we first run SSL pre-training with unlabeled data using PFML, after which we fine-tune our models for downstream classification tasks using labeled data. We compare PFML against three different baselines: MAE (He et al., 2022), data2vec (Baevski et al., 2022), and not using pre-training at all. We selected MAE for our experiments since it is conceptually very similar to PFML, and we chose data2vec since it is the current state-of-the-art data modality agnostic SSL method. In order to make the prediction of functionals directly comparable with predicting the input signal, we use a slightly modified version of MAE where we mask embeddings instead of masking inputs.
For PFML pre-training, our models consist of a modality-specific frame-level encoder (detailed in Sections 4.1, 4.2, and 4.3 for IMU, speech, and EEG data, respectively) and a Transformer network consisting of $T$ Transformer encoder blocks. Between the encoder and Transformer networks there is a CNN-based relative positional encoder followed by a GeLU (Hendrycks & Gimpel, 2016) activation and layer normalization (Ba et al., 2016). We frame our input signals before feeding the data into an encoder model, and we compute functionals from these frames as our training targets. For multi-channel data, we compute functionals separately for each channel. The functionals are then z-score normalized across the entire pre-training dataset. For computational efficiency, we pre-compute the functionals of each signal frame before the pre-training process. After the Transformer encoder blocks, we add a linear projection to convert the Transformer outputs into predicted functionals. Pre-training is run until validation loss convergence, and we use the model with the lowest validation loss as our pre-trained model. Starting from an initial learning rate, we gradually reduce the learning rate during model training with a reduction factor of 0.5 based on the plateauing of the validation loss.
@@
For pre-training and fine-tuning, we use the RAdam (Liu et al., 2020) and Adam (Kingma & Ba, 2015) optimizers, respectively. For the “no pre-training” condition, we simply omit pre-training, the first fine-tuning stage, and the learning rate warm-up period of the second fine-tuning stage. We used an NVIDIA Tesla V100 GPU to train our models, and we implemented the code using PyTorch version 1.13.1. Our implementation is publicly available on GitHub.222(Here will be a link to our GitHub repository.)
In order to demonstrate the superiority of PFML against the state-of-the-art SSL method for multiple data modalities, data2vec, in terms of representation collapse, we ran PFML, MAE, and data2vec pre-training 10 times using the best hyperparameter combinations for each SSL method and for each data modality. We defined representation collapse to have occurred if the variance of either the embeddings or model outputs fell below 0.01 for 10 consecutive pre-training epochs, during which the validation loss was decreasing. In our preliminary experiments, we found that this condition was a good indicator of an upcoming representation collapse: A systematic decrease in the variance of a model’s embeddings or outputs indicates impending representation collapse in SSL methods where the model can invent its own training targets.
@@ 4.3 Sleep Stage Classification
4.4 Results
Table 1: Downstream task fine-tuning results for PFML, data2vec, MAE, and not using pre-training at all for the five different classification tasks across the three different data modalities (IMU, speech, and EEG data).
@@ 4.4 Results
Table 1 presents the fine-tuning results of the comparison of our PFML method against MAE, data2vec, and not using pre-training at all. Across all three data modalities and five classification tasks, the results show that PFML outperformed MAE and achieved highly comparable results to data2vec. Using pre-training with any SSL method provided superior results as opposed to not using pre-training at all. For the classification of posture from IMU data, there were only minor differences in performance between different SSL methods. In sleep stage classification from EEG data, both MAE and PFML outperformed data2vec by a large margin. The comparison between PFML and MAE showcases that it is more beneficial to predict functionals than to predict the input signal.
Table 2: Linear evaluation results for PFML, data2vec, MAE, and a randomly initialized model.
Table 3: Frequency of representation collapse across 10 runs of PFML, data2vec, and MAE for each tested data modality.
@@
Table 2 shows the results of the linear evaluation experiments. Similar to the results of Table 1, PFML outperformed MAE and was comparable to data2vec when using the pre-trained models as feature extractors for linear classifiers. Again, both MAE and PFML outperformed data2vec by a large margin in sleep stage classification from EEG data. In the case of using a randomly initialized model as a feature extractor for linear classifiers, the classification accuracy was at chance-level in all cases except when classifying posture for IMU data.
@@
The results of representation collapse experiments are shown in Table 3. As can be seen from the results, it is very common for representation collapse to occur with data2vec across all data modalities. On the contrary, the results indicate that MAE and PFML do not suffer from representation collapse: PFML did not experience representation collapses at all, and MAE had a representation collapse only once. Furthermore, we attribute this single representation collapse of MAE to bad luck in model weight initialization, as in this particular case the model loss started diverging from the beginning of the pre-training process. The results showcase that methods like MAE and PFML, whose training targets inherently contain variance, are less prone to representation collapse compared to methods like data2vec that learn their own prediction targets.
4.5 Additional Hyperparameter Experiments

### v2

@@ Abstract
Self-supervised learning (SSL) is a data-driven learning approach that utilizes the innate structure of the data to guide the learning process. In contrast to supervised learning, which depends on external labels, SSL utilizes the inherent characteristics of the data to produce its own supervisory signal. However, one frequent issue with SSL methods is representation collapse, where the model outputs a constant input-invariant feature representation. This issue hinders the potential application of SSL methods to new data modalities, as trying to avoid representation collapse wastes researchers’ time and effort. This paper introduces a novel SSL algorithm for time-series data called Prediction of Functionals from Masked Latents (PFML). Instead of predicting masked input signals or their latent representations directly, PFML operates by predicting statistical functionals of the input signal corresponding to masked embeddings, given a sequence of unmasked embeddings. The algorithm is designed to avoid representation collapse, rendering it straightforwardly applicable to different time-series data domains, such as novel sensor modalities in clinical data. We demonstrate the effectiveness of PFML through complex, real-life classification tasks across three different data modalities: infant posture and movement classification from multi-sensor inertial measurement unit data, emotion recognition from speech data, and sleep stage classification from EEG data. The results show that PFML is superior to a conceptually similar SSL method and a contrastive learning-based SSL method. Additionally, PFML is on par with the current state-of-the-art SSL method, while also being conceptually simpler and without suffering from representation collapse. The code is freely available at https://github.com/SPEECHCOG/PFML.
@@ 1 Introduction
    We show that PFML obtains superior results against both a conceptually similar SSL method and a contrastive learning-based SSL method. Additionally, PFML achieves results that are on par with the current state-of-the-art data modality-agnostic SSL method, while also being conceptually simpler and without suffering from representation collapse.
@@ 2 Related Work
Similar to the advances in SSL for speech data, there have been significant developments in SSL for image data as well (Lee et al., 2017; Gidaris et al., 2018; Caron et al., 2018; Grill et al., 2020; Chen et al., 2020; Radford et al., 2021; He et al., 2022; Bao et al., 2022; Oquab et al., 2024). Grill et al. (2020) presented an SSL method that uses two neural networks that learn from each other’s representations of differently augmented views of the same image. He et al. (2022) proposed masked autoencoders (MAE) that try to reconstruct masked patches of input images using an asymmetric encoder-decoder architecture. The SSL algorithm by Bao et al. (2022) tokenizes images into visual tokens, followed by masking some image patches and then trying to recover the original tokens from the masked patches.
@@
SSL has also excelled in natural language processing (Devlin et al., 2019; Brown et al., 2020; Tay et al., 2023; OpenAI, 2023). Devlin et al. (2019) introduced an SSL method which obtains bidirectional feature representations from unlabeled text by conditioning on both the left and right textual context. The method by Brown et al. (2020) uses an autoregressive model which alternates dense and locally banded sparse attention patterns in their Transformer model. OpenAI (2023) proposed an expanded version of the method by Brown et al. (2020) by making the model not only larger, but also capable of handling image inputs in addition to text inputs.
@@
More recently, SSL literature has seen a growing number of work towards SSL algorithms capable of running the pre-training task on multiple different data modalities. The authors of van den Oord et al. (2018) developed an SSL approach that predicts future embeddings based on previous context using contrastive learning. They showed that their method was able to learn useful feature representations for audio, image, text, and reinforcement learning in 3D environments. The SSL method by Akbari et al. (2021) also uses contrastive learning, but their method simultaneously takes audio, video, and text data as input and creates multimodal feature representations. These features were shown to work well with multiple different downstream tasks, i.e. video action recognition, audio event classification, image classification, and text-to-video retrieval. Baevski et al. (2022) proposed data2vec, an SSL method for audio, image, and text data. In their approach, the model tries to predict masked latent features of an older version of itself that are both normalized and averaged over multiple Transformer layers. Their results in downstream tasks demonstrate the effectiveness of the method in all three data modalities. Yue et al. (2022) presented a universal framework, TS2Vec, for learning time-series representations at different semantic levels using hierarchical contrastive learning over augmented context views. Their approach enables robust contextual representations for each timestamp and allows for simple aggregation to obtain representations of arbitrary sub-sequences. They demonstrated that their method obtained state-of-the-art performance in time-series classification, forecasting, and anomaly detection tasks on multiple datasets, most of which were small in terms of the number of training samples. Notably, TS2Vec achieved these results by adding very simple classifiers after the pre-trained model, which has fixed constraints on the model architecture to model the hierarchical representations. Wang et al. (2023) proposed an SSL method that performs prediction of masked tokens in a unified manner on images, texts, and image-text pairs. Their experiments showed that their method achieves state-of-the-art performance on various vision and vision-language tasks.
@@
For modality-agnostic SSL algorithms, objective functions play a crucial role in guiding the learning process. These functions can be broadly categorized into three types: instance discrimination, clustering, and masked prediction. Instance discrimination aims to distinguish between different instances of data, thereby encouraging the model to learn unique features for each instance and enhancing the discriminative power of the learned representations. Contrastive learning methods, such as van den Oord et al. (2018); Baevski et al. (2020); He et al. (2020); Akbari et al. (2021); Pizzi et al. (2022); Yue et al. (2022), are an example of instance-discrimination-based SSL methods. Clustering, on the other hand, groups similar instances together in the feature space, fostering the model to learn common features among instances belonging to the same group. Methods like Caron et al. (2020); Asano et al. (2020); Hsu et al. (2021) are examples of clustering-based SSL methods. Lastly, masked prediction involves the task of predicting masked parts of the input data based on the unmasked parts, thereby encouraging the model to learn contextual relationships within the data. Examples of such SSL methods include Devlin et al. (2019); Wang et al. (2020); Baevski et al. (2022); He et al. (2022); Xie et al. (2022).
@@ 3.2 Prediction of Functionals from Masked Latents
Under these assumptions, as the model is trying to predict the computed functionals $\mathbf{f}_{n}$ given the embeddings $\mathbf{z}_{n}$, good model predictions $\mathbf{y}_{n}$ that lead to low prediction loss values also inherently contain variance. On the contrary, if $\mathbf{y}_{n}$ were to contain zero variance across the frames while $\mathbf{f}_{n}$ contains variance, the prediction loss would be high. Consequently, PFML pre-training does not converge to collapsed feature representations, as long as Assumptions 1 and 2 hold true. For a more detailed formulation, see Appendix A. Empirical results (see Section 5) support this theoretical claim, showing that PFML maintains variance in predictions across various datasets.
@@ 4 Experiments
We evaluate our PFML method using three different datasets of time-series data with complex classification tasks: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. For each dataset, we first run SSL pre-training with unlabeled data using PFML, after which we fine-tune our models for downstream classification tasks using labeled data. We compare PFML against four different baselines: MAE (He et al., 2022), data2vec (Baevski et al., 2022), TS2Vec (Yue et al., 2022), and not using pre-training at all. We selected MAE for our experiments since it is conceptually very similar to PFML, and we chose data2vec since it is the current state-of-the-art data-modality-agnostic SSL method. We also included the conceptually different TS2Vec in the present experiments due to its reported state-of-the-art performance on multiple different datasets of single- and multi-channel time-series data. In order to make the prediction of functionals directly comparable with predicting the input signal, we use a slightly modified version of MAE where we mask embeddings instead of masking inputs.
This section is organized as follows: First, Section 4.1 gives general-level information on the pre-training and fine-tuning experiments regarding PFML, MAE, and data2vec methods that utilize identical neural network architecture for the end-use system. This is followed by Sections 4.2, 4.3, and 4.4, which give modality-specific experimental details for IMU, speech, and EEG data, respectively. Finally, Section 4.5 explains the key differences between TS2Vec-related experiments and the experiments described in Section 4.1.
4.1 PFML, MAE, and data2vec Experiments
For PFML pre-training, our models consist of a modality-specific frame-level encoder (detailed in Sections 4.2, 4.3, and 4.4 for IMU, speech, and EEG data, respectively) and a Transformer network consisting of $T$ Transformer encoder blocks. Between the encoder and Transformer networks there is a CNN-based relative positional encoder followed by a GeLU (Hendrycks & Gimpel, 2016) activation and layer normalization (Ba et al., 2016). We frame our input signals before feeding the data into an encoder model, and we compute functionals from these frames as our training targets. For multi-channel data, we compute functionals separately for each channel. The functionals are then z-score normalized across the entire pre-training dataset. For computational efficiency, we pre-compute the functionals of each signal frame before the pre-training process. After the Transformer encoder blocks, we add a linear projection to convert the Transformer outputs into predicted functionals. After pre-training, this linear projection is discarded. Pre-training is run until validation loss convergence, and we use the model with the lowest validation loss as our pre-trained model. Starting from an initial learning rate, we gradually reduce the learning rate during model training with a reduction factor of 0.5 based on the plateauing of the validation loss.
@@ 4.1 PFML, MAE, and data2vec Experiments
For pre-training, we use the RAdam (Liu et al., 2020), AdamW (Loshchilov & Hutter, 2019), and Adam (Kingma & Ba, 2015) optimizers for PFML, MAE, and data2vec, respectively. For fine-tuning, we use the Adam optimizer. For the “no pre-training” condition, we simply omit pre-training, the first fine-tuning stage, and the learning rate warm-up period of the second fine-tuning stage. We used an NVIDIA Tesla V100 GPU to train our models, and we implemented the code using PyTorch version 1.13.1. Our implementation is publicly available on GitHub. [^2] https://github.com/SPEECHCOG/PFML
In order to demonstrate the superiority of PFML against the state-of-the-art SSL method for multiple data modalities, data2vec, in terms of representation collapse, we ran each of the pre-training algorithms of the present experiments 10 times using the best hyperparameter combinations for each SSL method and for each data modality. We defined representation collapse to have occurred if the variance of either the embeddings or model outputs fell below 0.01 for 10 consecutive pre-training epochs, during which the validation loss was decreasing. In our preliminary experiments, we found that this condition was a good indicator of an upcoming representation collapse: A systematic decrease in the variance of a model’s embeddings or outputs indicates impending representation collapse in SSL methods where the model can invent its own training targets.
@@ 4.4 Sleep Stage Classification
4.5 TS2Vec Experiments
We also conducted a similar set of experiments as described in Section 4.1 for the TS2Vec method. TS2Vec stands out from the other three SSL methods used in the present experiments due to its hierarchical contrastive learning algorithm, which is integrated into the pre-defined CNN encoder architecture. Therefore, TS2Vec does not allow the use of the same modality-specific encoders used with PFML, MAE, and data2vec in our experiments. Additionally, it is not possible to pre-train the classification model (i.e. Transformer) in TS2Vec. Consequently, for all three data modalities (IMU, speech, and EEG data), we use the pre-defined TS2Vec encoder, and we do not pre-train the Transformer network. In order to compare the linear separability of the SSL features, we add one linear layer directly after the TS2Vec pre-trained encoder, and we fine-tune this single linear layer while the weights of the encoder are frozen.
For TS2Vec pre-training, we used the original TS2Vec implementation [^3] https://github.com/zhihanyue/ts2vec with default hyperparameter settings, except for two modifications: First, we observed that the pre-defined number of minibatch updates was insufficient for our pre-training experiments. We found it more effective to run TS2Vec pre-training similarly to PFML, MAE, and data2vec, i.e. until validation loss convergence, and then use the model with the lowest validation loss as the final pre-trained model. Second, we adjusted the output dimensionality of the TS2Vec encoder to match the modality-specific input dimensionalities of the Transformer (see Table 5 of Appendix B).
For fine-tuning the TS2Vec pre-trained models, we added a randomly initialized Transformer model and two randomly initialized fully-connected GeLU layers followed by a softmax function after the TS2Vec pre-trained encoder. The models were then fine-tuned similarly to the process described in Section 4.1, except that in the first fine-tuning stage, both the Transformer and the GeLU layers were fine-tuned while the weights of the encoder were frozen. The fine-tuning hyperparameters were the same as those used for the other pre-training algorithms (see Table 6 of Appendix B).
To demonstrate fair comparison, we also conducted three additional experiments using TS2Vec: First, similar to Section 4.1, we tested the “no pre-training” condition of the neural network architecture combining the TS2Vec encoder and Transformer classifier by fine-tuning the entire model at once using a randomly initialized model. Second, we tested the linear separability of a randomly initialized TS2Vec encoder by adding a single linear layer after the encoder and fine-tuning only this layer. Third, since the authors of the TS2Vec paper (Yue et al., 2022) used an SVM classifier with an RBF kernel for the SSL features in classification tasks, we adopted the same procedure for our experimental pipeline. All of the TS2Vec-related results are presented in Appendix F.
5 Results
@@
Table 1 presents the fine-tuning results of the comparison of our PFML method against MAE, data2vec, TS2Vec, and not using pre-training at all. Across all three data modalities and five classification tasks, the end-use results show that PFML outperformed MAE and TS2Vec, and achieved results that are on par with data2vec. Using pre-training with any SSL method other than TS2Vec provided superior results as opposed to not using pre-training at all. Furthermore, for the classification of posture from IMU data, which is considered an easy task (Airaksinen et al., 2022), there were only minor differences in performance between all SSL methods other than TS2Vec. The comparably weak results for TS2Vec are most probably due to the fact that TS2Vec cannot utilize modality-specific encoders, and also since the Transformer model is not pre-trained during TS2Vec pre-training. Using SVM models instead of Transformers in TS2Vec classification, as in the original article of Yue et al. (2022), did not improve the results (see Appendix F). In sleep stage classification from EEG data, both MAE and PFML outperformed data2vec by a large margin. Also, the comparison between PFML and MAE showcases that it is more beneficial to predict functionals than to predict the input signal.
@@
Table 2 shows the results of the linear evaluation experiments. Similar to the results of Table 1, PFML outperformed MAE and was comparable to data2vec when using the pre-trained models as feature extractors for linear classifiers. Again, both MAE and PFML outperformed data2vec by a large margin in sleep stage classification from EEG data. In valence classification from speech data and sleep stage classification from EEG data, the TS2Vec method excelled. However, for other classification tasks, the SSL features provided by TS2Vec were the weakest performance-wise. In the case of using a randomly initialized model as a feature extractor for linear classifiers, the classification accuracy was at chance-level in all cases except when classifying posture for IMU data.
Table 1: Downstream task fine-tuning results for PFML, data2vec, MAE, TS2Vec, and not using pre-training at all for the five different classification tasks across the three different data modalities (IMU, speech, and EEG data).
44footnotetext: TS2Vec uses a different encoder compared to other experiments, since the TS2Vec algorithm is built into the pre-defined CNN encoder architecture.
Table 2: Linear evaluation results for PFML, data2vec, MAE, TS2Vec, and a randomly initialized model.
55footnotetext: In TS2Vec, the linear layer is added directly after the pre-trained encoder.
Table 3: Frequency of representation collapse across 10 runs of PFML, data2vec, MAE, and TS2Vec for each tested data modality.
@@
The results of representation collapse experiments are shown in Table 3. As can be seen from the results, it is very common for representation collapse to occur with data2vec across all data modalities. On the contrary, the results indicate that PFML, MAE, and TS2Vec do not suffer from representation collapse: PFML and TS2Vec did not experience representation collapses at all, and MAE had a representation collapse only once. Furthermore, we attribute this single representation collapse of MAE to bad luck in model weight initialization, as in this particular case the model loss started diverging from the beginning of the pre-training process. The results showcase that methods like PFML, MAE, and TS2Vec, whose training targets inherently contain variance, are less prone to representation collapse compared to methods like data2vec that learn their own prediction targets.
6 Additional Hyperparameter Experiments